OpenStudy (anonymous):

Simplify: 4 square root of 2 end root plus 7 square root of 2 end root minus 3 square root of 2

OpenStudy (luigi0210):

Is it this? \(\large 4\sqrt{2}+7\sqrt{2}-3\sqrt{2}\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Sorry about the wait >.< But all you have to do is add them together

OpenStudy (anonymous):

i did that didnt give me the right answer

OpenStudy (luigi0210):

Did you get \(8\sqrt{2}\)?

OpenStudy (anonymous):

no here the 2 is i got 6 lol idont know why

OpenStudy (anonymous):

can you check my work?i have 8 problems left

OpenStudy (luigi0210):

Just open new questions, I have to leave soon anyways, sorry :/